OpenStudy (anonymous):

A 16 foot ladder is leaning against a wall. If the top slips down the wall at a rate of 3 ft/s, how fast will the foot be moving away from the wall when the top is 13 feet above the ground?

OpenStudy (anonymous):

I Think i know how to do this but its not accepting my answer

OpenStudy (anonymous):

I drew my triangle 16ft 13ft 9.327ft

OpenStudy (anonymous):

then i took the a^2+b^2=c^2 and took the derivative

OpenStudy (anonymous):

i plugged in my numbers for a b c then plugged in for 1 for the rate of c and -3 for the rate of b

OpenStudy (ranga):



OpenStudy (anonymous):

y = 13 correct?

OpenStudy (anonymous):

x should be 9.27

OpenStudy (ranga):

Before you put in 13, you have to leave it as the variable y. Use x^2 + y^2 = 16^2 Differentiate throughout with respect to t. Tell me what you get.

OpenStudy (anonymous):

2x(dx/dt)+2y(dy/dt)=256

OpenStudy (anonymous):

well actually = 0

OpenStudy (ranga):

right. Now put dy/dt = -3, y = 13. Find x using pythagoras theorem and put that value and solve for dx/dt.

OpenStudy (anonymous):

2x(dx/dt)-78=256

OpenStudy (anonymous):

so 2x(dx/dt)=334

OpenStudy (ranga):

there is no 256. The right side is 0 when you take the derivative.

OpenStudy (anonymous):

shoot sorry i forgot 2x(dx/dt) = 78

OpenStudy (ranga):

2x(dx/dt) + 2y(dy/dt) = 0 Divide by 2 throughout x(dx/dt) + y(dy/dt) = 0 dy/dt = -3, y = 13 x(dx/dt) = 39 dx/dt = 39/x. From the right angle we know x = sqrt(16^2 - 13^2) Solve for dx/dt

OpenStudy (anonymous):

ahhhh i see my problem

OpenStudy (anonymous):

i was taking the derivative first instead of = 0

OpenStudy (anonymous):

i got 4.1812 for my final

OpenStudy (ranga):

cool. I don't know what decimal accuracy it needs. But 4.18 or 4.2 should be acceptable..

OpenStudy (anonymous):

just tried it. it worked. Thanks for the help. When doing problems with the pathagoreon theorem like these do you always make your C 0?

OpenStudy (ranga):

Not sure what you mean by "make you C 0".

OpenStudy (anonymous):

well i mean do you always plug in C before taking the derivative?

OpenStudy (ranga):

C is the ladder length and it is fixed at 16 feet. So no matter how low or high the ladder rests, the x^2 and y^2 will always add up to the square of the ladder length. So x^2 + y^2 = 16^2.

OpenStudy (anonymous):

ohhh ok i understand. so because the ladder is a constant

OpenStudy (ranga):

But when you take the derivative, the right hand side will be zero because it is a constant.

OpenStudy (anonymous):

I understand. makes alot more sense now

OpenStudy (ranga):

cool.